# Probabilistic estimates of the diameters of the Rubik’s Cube groups

### v5

@@ VI.2 Quarter-turn metric
meaning that the predicted diameter of 48 is definitive, which however does not imply the accuracy of the prediction.
@@ VIII Discussion
Table 11 summarizes the size and metric dependence of the actual (if available) and predicted diameters of the $n\times n\times n$ Rubik’s Cube groups ($2\leq n\leq 5$). This table’s predictions can be easily extended to greater $n$, different metrics, or other types of Cubes with essentially little additional computational cost, but determining actual values will be formidable.
@@
Cube Metric $r$ [^1] Branching ratio. See Eq. (4). Actual Predicted [^2] Probabilistic estimate based on a modified version of the coupon collector’s problem (this work). Predicted [^3] Equation (101) unless otherwise noted (this work).
2$\times$2$\times$2 Half 5.94 11444Reference Ferenc, 2024. 12 11
2$\times$2$\times$2 Quarter 4.44 14444Reference Ferenc, 2024. 14 14
2$\times$2$\times$2 Semi-quarter 2.77 19555This work. 21 20
2$\times$2$\times$2 Bi-quarter 9.19 10555This work. 9 9
3$\times$3$\times$3 Half 13.33 20666Rokicki et al.Rokicki et al. (2014b, 2013, a) 22 21
3$\times$3$\times$3 Quarter 9.37 26666Rokicki et al.Rokicki et al. (2014b, 2013, a) 26 25
4$\times$4$\times$4 Half 20.67 $\cdots$ 41 40
4$\times$4$\times$4 Quarter 14.30 $\cdots$ 48 47
5$\times$5$\times$5 Half 28.08 $\cdots$ 58 58
5$\times$5$\times$5 Quarter 19.23 $\cdots$ 68 67
@@ VIII.1 Comparison between predicted and actual results
For the 2$\times$2$\times$2 Cube, one can easily perform the brute-force enumeration of configurations as a function of the number of turns applied to the solved configuration, in nearly but not exactly the same way the probabilistic estimation’s hypothetical algorithm generates configurations.
In Figs. 14 are plotted the number of new, non-redundant configurations generated (“Actual”) in each step for the half-turn, quarter-turn, semi-quarter-turn, and bi-quarter-turn metrics, respectively, which are compared with the probabilistic estimates (“Predicted”). The “Actual” numbers were determined by a computer program.Hirata (2014) The “Predicted” numbers were obtained as
@@
Implicit in our probabilistic estimation logic is the assumption that the details of each metric are unimportant and all of its characteristics can be encapsulated into the branching ratio $r$, which is invariant of the step count. From Figs. 14, this assumption seems valid for the growth phase, but is no longer so in the decay phase. The fact that these curves show a super-exponential decay suggests that the effective value of $r$ increases near the diameter and may even exceed the number of turns. In fact, the actual and predicted curves in Fig. 3 behave like $e^{-4.0t}$ and $e^{-2.6t}$(where $t$ is the number of steps) at the diameter, respectively, and the factor of 4.0 in the former exceeds $r=2.77$ or even the number of turns, which is three. There seems room for improvement in the probabilistic estimates in the peak position and decay rate of these curves.
@@
Figure 5: The number of new (non-redundant) configurations of the 3$\times$3$\times$3 Cube generated in each step in the half-turn metric. The actual values (those in the last five steps were approximate) were taken from Table 5.1 of Rokicki et al.Rokicki et al. (2013)
@@
For the 3$\times$3$\times$3 Cube, Rokicki et al.Rokicki et al. (2013) gave the actual and estimated numbers (“positions”) of new configurations generated at each step. They are compared with our predicted values in Fig. 5. In this case, the rising part and the peak of the actual curve is accurately reproduced by the prediction, but the falling part decays too slow in the latter, overestimating the diameter by two, although the Rokicki et al.’s values in the last five steps are approximate. The actual curve has a rapid falloff of $e^{-22.3t}$, although it likely decreases faster than exponentially. On the other hand, the predicted curve decays exponentially as $e^{-13.3t}$ corresponding to $r=13.33$. It again suggests that our probabilistic logic tends to underestimate the accelerated rate of exhausting remaining unrealized configurations in the final few steps. However, the exact agreement of the predicted and actual diameters in the quarter-turn metric may contradict this view, given the smaller value of $r=9.37$ and even slower decay of the predicted curve in this metric.
@@
Singmaster,Singmaster (1981) using the same argument, inferred that nearly all configurations should be generated at 17.3 steps, which is only slightly less than our value or the one of Rokicki et al.Rokicki et al. (2013) On this basis, Singmaster conjectured that “every position can be achieved in at most 20 moves,” which is in exact agreement with the correct value determined by Rokicki et al.Rokicki et al. (2013) some 33 years later, although no account was given about the addition of 3 moves.
@@
In the 4$\times$4$\times$4 and 5$\times$5$\times$5 Cubes, whose branching ratio $r$ is small relative to the huge $E[T_{N}]$, there expected to be long decay phases, potentially causing noticeable overestimation of the diameters.
@@ VIII.2 Size dependence
Demaine et al.Demaine et al. (2011) argued that the diameter is asymptotically proportional to $n^{2}/\log n$, where $n$ denotes the Cube size. Under the assumption of the exact proportionality in the quarter-turn metric, their formula predicts the diameters of the 4$\times$4$\times$4 and 5$\times$5$\times$5 Cubes in the same metric to be 43 and 62, respectively, which are much smaller than our estimates (48 and 68).
Our probabilistic estimation suggests that a more meaningful closed-form expression may be a function of the number of configurations $N$ (rather than of the Cube size $n$) and also of the metric. This is because there are complicated $n$ dependencies (possibly including even-odd alternation) of $N$ and also because it is $N$ that dictates the complexity of the puzzle and thus the diameters more directly than $n$. Furthermore, as discussed below, the diameters may also depend on the metric in a nontrivial manner since some turns seem more useful than others.
@@
The first term counts the number of steps required for $S(t)$ to increase geometrically and reach $N$, while the second term is the number of additional steps needed for $T(t)$ to increase linearly and exceed $E[T_{N}]\approx N\ln N$, in both cases approximately. This formula gives the predicted value of the diameter within one from either the correct or probabilistically predicted value for any given Cube size $n$ and metric.
@@ Acknowledgements.
Suppose that $m$ (out of total $N$) non-redundant configurations have been generated so far. How many ($t_{m}$) random configurations need to be generated before one more non-redundant configuration is added? Let $E[t_{m}]$ be the expectation value of this number. The probability that the first randomly generated configuration is non-redundant is
$p_{m}=\frac{N-m}{N}.$ (102)
The probability that the $k$th randomly generated configuration is non-redundant for the first time is
$(1-p_{m})^{k-1}p_{m}.$ (103)
Writing $q_{m}=1-p_{m}$, the expected value $E[t_{m}]$ is then an infinite sum,
$E[t_{m}]$ $=$ $1\cdot p_{m}+2\cdot q_{m}p_{m}+3\cdot q_{m}^{2}p_{m}+\dots$ (104)
$=$ $p_{m}\frac{d}{dq_{m}}\left(q_{m}+q_{m}^{2}+q_{m}^{3}+\dots\right)$
$=$ $p_{m}\frac{d}{dq_{m}}\frac{q_{m}}{1-q_{m}}$
$=$ $p_{m}\frac{1}{(1-q_{m})^{2}}=\frac{1}{p_{m}},$
where we used $\lim_{k\to\infty}q_{m}^{k}=0$. Variance $V[t_{m}]$ is also evaluated similarly as
$V[t_{m}]$ $=$ $\Big{(}E[t_{m}^{2}]\Big{)}-\Big{(}E[t_{m}]\Big{)}^{2}$ (105)
$=$ $\left(1^{2}\cdot p_{m}+2^{2}\cdot q_{m}p_{m}+3^{2}\cdot q_{m}^{2}p_{m}+\dots\right)-\frac{1}{p_{m}^{2}}$
$=$ $p_{m}\frac{d}{dq_{m}}\left(q_{m}+2q_{m}^{2}+3q_{m}^{3}+\dots\right)-\frac{1}{p_{m}^{2}}$
$=$ $p_{m}\frac{d}{dq_{m}}q_{m}\frac{d}{dq_{m}}\left(q_{m}+q_{m}^{2}+q_{m}^{3}+\dots\right)-\frac{1}{p_{m}^{2}}$
$=$ $p_{m}\frac{d}{dq_{m}}q_{m}\frac{d}{dq_{m}}\frac{q_{m}}{1-q_{m}}-\frac{1}{p_{m}^{2}}$
$=$ $p_{m}\frac{d}{dq_{m}}\frac{q_{m}}{(1-q_{m})^{2}}-\frac{1}{p_{m}^{2}}$
$=$ $\frac{1}{p_{m}^{2}}-\frac{1}{p_{m}}.$
Let $E[T_{N}]$ be the expected number of random configurations that need to be generated so that every one of $N$ configurations occurs at least once. It should be the sum of $E[t_{m}]$ over $10\leq m\leq N-$ 1.
@@
$=$ $N^{2}\left(\frac{1}{1^{2}}+\frac{1}{2^{2}}+\frac{1}{3^{2}}+\dots+\frac{1}{N^{2}}\right)-E[T_{N}]$
@@
$\sum_{n=1}^{\infty}\frac{1}{n^{2}}=\frac{\pi^{2}}{6}.$ (108)
@@
What is the expected number $E[N_{T}]$ of non-redundant configurations among $T$ randomly generated configurations, when the total number of possible configurations is $N$? Let $p_{t}$ be the probability that the $t$th randomly generated configuration is distinct from all previously generated configurations. Furthermore, let $N_{t}$ be the number of non-redundant configurations after $t$ random configurations are generated. We have
$p_{t+1}=\frac{N-N_{t}}{N},$ (110)
@@
$N_{t}=p_{t}\left(N_{t-1}+1\right)+\left(1-p_{t}\right)N_{t-1}=N_{t-1}+p_{t}.$ (112)
@@
$p_{t+1}$ $=$ $\left(1-\frac{1}{N}\right)p_{t}=\left(1-\frac{1}{N}\right)^{t},$ (113)
where we used $p_{1}=1$ and $N_{0}=0$. Therefore, after $T$ random configurations are generated, the expected number of only the non-redundant configurations among all the generated configurations is
$E[N_{T}]$ $=$ $N(1-p_{T+1})$ (114)
$=$ $N\left\{1-\left(1-\frac{1}{N}\right)^{T}\right\}$
$\approx$ $N\left(1-e^{-T/{N}}\right),$
in the large $N$ limit.
For $T\ll N$, approximating the exponential by its first-order Taylor expansion, we have
$E[N_{T}]=N\left(1-e^{-T/{N}}\right)\approx N\left\{1-\left(1-\frac{T}{N}\right)\right\}=T,$ (115)
indicating that most of all $T$ generated configurations are non-redundant.
The expected number of configurations that are not realized even once is
$N-E[N_{T}]=N-N\left(1-e^{-T/{N}}\right)=Ne^{-T/{N}}.$ (116)
which is consistent with the Siobhan distribution [Eq. (3)]. When this number is smaller than one, it is expected that all of the $N$ configurations have been realized (the coupon collector’s problem).
$Ne^{-T/N}<1,$ (117)
@@
$N\ln N<T.$ (118)
This is consistent with (if not identical to) the solution of the coupon collector’s problem given in Appendix A.

### v6

@@ VI.2 Quarter-turn metric
meaning that the predicted diameter of 48 is definitive, which however says nothing about the accuracy of the prediction.
@@ VIII Discussion
Table 12 summarizes the size and metric dependence of the actual (if available) and predicted diameters of the $n\times n\times n$ Rubik’s Cube groups ($2\leq n\leq 5$). This table’s predictions can be easily extended to greater $n$, different metrics, or other types of Cubes with essentially little additional computational cost, but determining actual values will be formidable in many cases.
@@
Cube Metric $r$ [^1] Branching ratio. See Eq. (4). Actual Predicted [^2] Probabilistic estimate based on a modified version of the coupon collector’s problem (this work). Predicted [^3] Equation (111) (this work).
2$\times$2$\times$2 Half 5.94 11444Reference Ferenc (2024). 12 11.0
2$\times$2$\times$2 Quarter 4.44 14444Reference Ferenc (2024). 14 13.5
2$\times$2$\times$2 Semi-quarter 2.77 19555This work. 21 20.3
2$\times$2$\times$2 Bi-quarter 9.19 10555This work. 9 8.5
3$\times$3$\times$3 Half 13.33 20666Rokicki et al. Rokicki et al. (2024, 2013, 2014). 22 20.8
3$\times$3$\times$3 Quarter 9.37 26666Rokicki et al. Rokicki et al. (2024, 2013, 2014). 26 25.0
3$\times$3$\times$3 Square 4.44 15555This work. 13 12.0
4$\times$4$\times$4 Half 20.67 $\cdots$ 41 40.0
4$\times$4$\times$4 Quarter 14.30 $\cdots$ 48 47.1
5$\times$5$\times$5 Half 28.08 $\cdots$ 58 57.5
5$\times$5$\times$5 Quarter 19.23 $\cdots$ 68 66.9
@@ VIII.1 Comparison between predicted and actual results
For the 2$\times$2$\times$2 Cube, one can easily perform the brute-force enumeration of configurations as a function of the number of turns applied to the solved configuration, in nearly but not exactly the same way in which the probabilistic estimation’s hypothetical algorithm generates configurations.
In Figs. 14 are plotted the number of new, non-redundant configurations generated (“Actual”) in each step for the half-turn, quarter-turn, semi-quarter-turn, and bi-quarter-turn metrics, respectively, which are compared with the probabilistic estimates (“Predicted”). The “Actual” numbers were determined by a computer program Hirata (2024a). The “Predicted” numbers were obtained as
@@
Implicit in our probabilistic estimation logic is the assumption that the details of each metric are unimportant and all of its characteristics can be encapsulated into the branching ratio $r$, which is invariant of the step count. From Figs. 14, this assumption seems largely valid for the growth phase, but is no longer so in the decay phase.
@@
Figure 5: The number of new (non-redundant) configurations of the 3$\times$3$\times$3 Cube generated in each step in the square-turn metric.
The 3$\times$3$\times$3 Cube in the square-turn metric can be solved by brute force because its number of configurations is relatively small ($N=6.63\times 10^{5}$). This is an interesting case: It is the same as the 2$\times$2$\times$2 Cube in the quarter-turn metric in every respect except for $N$, but the probabilistic estimate works most poorly for the former, while it gives an exact prediction for the latter. In fact, the diameter for the 3$\times$3$\times$3 Cube group in the square-turn metric with a smaller $N$ is greater by one than the diameter for the 2$\times$2$\times$2 Cube group in the quarter-turn metric with a larger $N$; the order and diameter of locally isomorphic graphs are inverted Hirata (2024b).
Figure 5 attests to the large deviations in the actual and predicted numbers of new configurations for the 3$\times$3$\times$3 Cube in the square-turn metric. The errors are seen already around the peaks and aggravated in the falling part. Unlike the foregoing cases, the decay of the actual curve is unusually slow, resulting in the inversion, and slower than the predicted curve’s decay. The cause of the slow decay is unknown, but the poor performance of the probabilistic logic may be ascribed to the relative smallness of $N$.
Figure 6: The number of new (non-redundant) configurations of the 3$\times$3$\times$3 Cube generated in each step in the half-turn metric. The actual values (those in the last five steps were approximate) were taken from Rokicki et al. Rokicki et al. (2013, 2024).
Figure 7: The number of new (non-redundant) configurations of the 3$\times$3$\times$3 Cube generated in each step in the quarter-turn metric. The actual values (those in steps 19 through 26 were approximate) were taken from Rokicki et al. Rokicki et al. (2024).
For the 3$\times$3$\times$3 Cube in the half- and quarter-turn metrics, Rokicki et al. Rokicki et al. (2013, 2024) gave the actual and estimated numbers (“positions”) of new configurations generated at each step. They are compared with our predicted values in Figs. 6 and 7.
@@
For the half-turn metric (Fig. 6), the rising part and the peak of the actual curve is accurately reproduced by the prediction, but the falling part decays too slow in the latter, overestimating the diameter by two. The actual curve has a rapid falloff of $e^{-22.3t}$, and it likely decreases even faster than exponentially. On the other hand, the predicted curve decays exponentially as $e^{-13.3t}$ corresponding to $r=13.33$. It again suggests that our probabilistic logic tends to underestimate the accelerated rate of exhausting remaining unrealized configurations in the final few steps.
@@
Singmaster Singmaster (1981), using the same argument, inferred that nearly all configurations should be generated at 17.3 steps, which is only slightly less than our value or the one of Rokicki et al. Rokicki et al. (2013). On this basis, Singmaster speculated that “every position can be achieved in at most 20 moves,” which is in exact agreement with the correct value determined by Rokicki et al. Rokicki et al. (2013) long before the latter became available. No account seems to have been given about the addition of the 3 moves, however.
For the quarter-turn metric (Fig. 7), on the other hand, the rising part of the actual curve is accurately reproduced by the prediction, while the falling part of the predicted curve initially underestimates and then overestimates the rate of decay. This is simply because the “actual” curve (which is only approximate Rokicki et al. (2024) for steps 19 through 26) shows an unusual (as compared with all the other cases discussed above) or even puzzling behavior of slowing down the rate of decay in the last three steps. The Cayley graph of this metric is expected to be a symmetric graph because every configuration is equivalent and so is every turn Hirata (2024b). For such a graph, the branching ratio $r$ is a monotonically decreasing function of step count Hirata (2024b), which seems to be contradicted by the “actual” curve Rokicki et al. (2024).
Overall, however, the agreement is good in Figs. 6 and 7, supporting the notion that probabilistic logic generally works progressively well for larger $N$.
@@
In the 4$\times$4$\times$4 and 5$\times$5$\times$5 Cubes, whose branching ratio $r$ is small relative to the huge $E[T_{N}]$, there expected to be long decay phases, potentially causing noticeable overestimation of the diameters. At the same time, it is also possible that their huge $N$ make the probabilistic estimates more accurate.
@@ VIII.2 Size dependence
Demaine et al. Demaine et al. (2011) argued that the diameter is asymptotically proportional to $n^{2}/\log n$ for the $n\times n\times n$ Cube. Under the assumption of the exact proportionality in the quarter-turn metric, their formula predicts the diameters of the 4$\times$4$\times$4 and 5$\times$5$\times$5 Cubes in the same metric to be 43 and 62, respectively, which are much smaller than our estimates (48 and 68).
Our probabilistic estimation suggests that a more meaningful closed-form expression may be a function of the number of configurations $N$ (rather than of the Cube size $n$) and also of the metric. This is because there are complicated $n$ dependencies (possibly including even-odd alternation) of $N$ and also because it is $N$ that dictates the complexity of the puzzle and thus the diameters more directly than $n$. Furthermore, the diameters may also depend on the metric in a nontrivial manner since some turns may be more useful than others.
@@
The first term counts the number of steps required for $S(t)$ to increase geometrically and reach $N$, while the second term is the number of additional steps needed for $T(t)$ to increase linearly and exceed $E[T_{N}]\approx N\ln N$, in both cases approximately. This formula gives the predicted value of the diameter within one from either the correct or probabilistically predicted value except for the smallest case (the 3$\times$3$\times$3 Cube in the square-turn metric), where the error is three.
In accompanying article Hirata (2024b), a lower bound for the diameter of symmetric Cayley graphs of some Rubik’s Cube groups is obtained on the basis of graph-theoretical arguments, and compared with lower and upper bounds for the diameter of random regular graphs of Bollobás and de la Vega Bollobás and de la Vega (1982); Bollobás (2001), who combined graph-theoretical and probabilistic arguments.
@@ Acknowledgements.
Let us consider a process of uniformly randomly generating an integer in the range of 1 through $N$. How many times ($T_{N}$) does the random generation need to be repeated before every integer in the range of 1 through $N$ occurs at least once? Let $E[T_{N}]$ be the expectation value of this number.
First, suppose that $n$ distinct integers have appeared so far after some number of random generations. How many additional times ($t_{n}$) does one need to carry out the random generations before a new, distinct integer is added? Let $E[t_{n}]$ be the expectation value of this number.
The probability that the first randomly generated integer is distinct is
$p_{n}=\frac{N-n}{N}.$ (112)
The probability that the first randomly generated integer is not distinct, but the second generated integer is is
$(1-p_{n})p_{n}.$ (113)
Generally, the probability that the $k$th random generation produces a new, distinct integer for the first time is
$(1-p_{n})^{k-1}p_{n}.$ (114)
Writing $q_{n}=1-p_{n}$, the expectation value $E[t_{n}]$ is then an infinite sum,
$E[t_{n}]$ $=$ $1\cdot p_{n}+2\cdot q_{n}p_{n}+3\cdot q_{n}^{2}p_{n}+\dots$ (115)
$=$ $p_{n}\frac{d}{dq_{n}}\left(q_{n}+q_{n}^{2}+q_{n}^{3}+\dots\right)$
$=$ $p_{n}\frac{d}{dq_{n}}\frac{q_{n}}{1-q_{n}}$
$=$ $p_{n}\frac{1}{(1-q_{n})^{2}}$
$=$ $\frac{1}{p_{n}}=\frac{N}{N-n}.$
The corresponding variance $V[t_{n}]$ is
$V[t_{n}]$ $=$ $\Big{(}E[t_{n}^{2}]\Big{)}-\Big{(}E[t_{n}]\Big{)}^{2}$ (116)
$=$ $\left(1^{2}\cdot p_{n}+2^{2}\cdot q_{n}p_{n}+3^{2}\cdot q_{n}^{2}p_{n}+\dots\right)-\frac{1}{p_{n}^{2}}$
$=$ $p_{n}\frac{d}{dq_{n}}\left(q_{n}+2q_{n}^{2}+3q_{n}^{3}+\dots\right)-\frac{1}{p_{n}^{2}}$
$=$ $p_{n}\frac{d}{dq_{n}}q_{n}\frac{d}{dq_{n}}\left(q_{n}+q_{n}^{2}+q_{n}^{3}+\dots\right)-\frac{1}{p_{n}^{2}}$
$=$ $p_{n}\frac{d}{dq_{n}}q_{n}\frac{d}{dq_{n}}\frac{q_{n}}{1-q_{n}}-\frac{1}{p_{n}^{2}}$
$=$ $p_{n}\frac{d}{dq_{n}}\frac{q_{n}}{(1-q_{n})^{2}}-\frac{1}{p_{n}^{2}}$
$=$ $\frac{1}{p_{n}^{2}}-\frac{1}{p_{n}}=\frac{N^{2}}{(N-n)^{2}}-E[t_{n}].$
Then, $E[T_{N}]$ is the sum of the average number $E[t_{n}]$ of random generation needed to produce the ($n$+1)th distinct integer over $10\leq n\leq N-$ 1. Therefore, in the large-$N$ limit,
@@
$=$ $N^{2}\left(\frac{1}{N^{2}}+\frac{1}{(N-1)^{2}}+\dots+\frac{1}{2^{2}}+\frac{1}{1^{2}}\right)-E[T_{N}]$
@@
$\sum_{k=1}^{\infty}\frac{1}{k^{2}}=\frac{\pi^{2}}{6}.$ (119)
@@
Consider a process of uniformly randomly generating $n$ integers in the range of 1 through $N$. How many ($N_{n}$) distinct integers are expected to be present in this set of $n$ integers? Let $E[N_{n}]$ be this expectation value.
Let $p_{k}$ be the probability that the $k$th randomly generated integer is distinct from all previously generated ones. Let $N_{k}$ be the number of distinct integers after the random generation is repeated $k$ times. We have
$p_{k+1}=\frac{N-N_{k}}{N},$ (121)
@@
$N_{k}=p_{k}\left(N_{k-1}+1\right)+\left(1-p_{k}\right)N_{k-1}=N_{k-1}+p_{k}.$ (123)
@@
$p_{k+1}$ $=$ $\left(1-\frac{1}{N}\right)p_{k}=\left(1-\frac{1}{N}\right)^{k},$ (124)
where we used $p_{1}=1$ and $N_{0}=0$. Therefore, after $n$ random integers are generated, the expected number of distinct integers among them is
$E[N_{n}]$ $=$ $N(1-p_{n+1})$ (125)
$=$ $N\left\{1-\left(1-\frac{1}{N}\right)^{n}\right\}$
$\approx$ $N\left(1-e^{-n/{N}}\right),$
in the large-$N$ limit.
For $n\ll N$, approximating the exponential by its first-order Taylor expansion, we have
$E[N_{n}]=N\left(1-e^{-n/{N}}\right)\approx N\left\{1-\left(1-\frac{n}{N}\right)\right\}=n,$ (126)
indicating that virtually all of $n$ randomly generated integers are distinct.
The expected number of integers that are not generated even once after $n$ random generations is
$N-E[N_{n}]=N-N\left(1-e^{-n/{N}}\right)=Ne^{-n/{N}}.$ (127)
When this number is smaller than one, it is expected that all of the $N$ integers have been realized.
$Ne^{-n/N}<1,$ (128)
@@
$N\ln N<n.$ (129)
This is consistent with the solution of the coupon collector’s problem given in Appendix A.